Fermi National Laboratory


Accelerator Update

Monday April 5
The day shift began with Tevatron (TeV) in shot setup, with the Antiproton Source (Pbar) and Recycler stashing antiprotons, and with beam to MiniBooNE.

Pbar resumed stacking at 8:35 AM.

At 8:43 AM, Operations established store 3359 with a luminosity of 48.36E30.

Operation prepared the Main Injector (MI) at 2:04 PM for an access and for the NuMI magnet testing.

At 2:14 PM, Operations began preparing Linac and the Preaccelerator (Preacc) for access.

Linac and the Preacc began tuning beam at 4:03 PM.

The MI access for NuMI magnet testing ended at 4:38 PM, but then other technicians accessed MI to check on Recycler vacuum issues. They came out about an hour later.

Operations established beam to MiniBooNE at 6:04 PM and stacking at 8:16 PM.

Operations began sending beam to Switchyard and the Meson MCenter experiment around 7:30 PM.

At 9:34 PM, two Booster RF stations (BRF5 and 10) tripped off and wouldn't reset. Operators called in a RF tech and he had the stations operating by 11:27 PM.

Tuesday April 6
The midnight shift began with Operations monitoring store 3359, with Pbar stacking, with Meson and MiniBooNE taking beam, and with Booster collimator experts investigating the beam tune.

At 1:37 AM, operators noticed the stack's core go coherent every once in a while, but couldn't understand why. They thought the problem may have been tied to an Accumulator RF station (ARF1).

A MI expert began a slow acceleration study at 5:52 AM.

A NuMI expert began a study period at 6:40 AM.

Operations shut off the beam to MCenter at 9:30 AM so the experimenters could access their collision hall.

At 1 PM, Operations terminated store 3359. By 1:13 PM operators had the Collider Detector at Fermilab (CDF) and D-zero (D0) collision halls ready for access.

A number of Pbar amplifiers (TWTs) that control the Accumulator's stacktail momentum tripped off at 2:41 PM and kept tripping off. An expert was informed and he began an investigation.

Operations put the TeV into shot setup at 3:14 PM and at 6 PM established store 3361 with an initial luminosity of 43.50E30. Pbar resumed stacking.

At 7:11 PM, the Recycler began stashing antiprotons from Pbar. Pbar had some closure problems, but by 7:37 PM, the Recycler had stashed about 17.5 mA of the 25 mA of antiprotons sent.

Operations reestablished beam to SY and MCenter at 9:19 PM.

Wednesday April 7
The midnight shift began with Operations monitoring store 3361, with Pbar stacking, with the Recycler maintaining its stash, and with MCenter and MiniBooNE taking beam.

At 4:51 AM, a Booster kicker power supply (B:MKS02) tripped off. Operators tried to reset the supply, but heard arcing from the cabinet and so left the supply off. Operations will notify the proper people in the morning.

Today's Plans
The plan for today is to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/7/2004   email Fermilab